OpenStudy (anonymous):

PLEASE help! FAN + MEDAL: Which value is equivalent to (8x4x2/8x7)^2 x (8^0/7^-3)^3 x 7^-9? 64/49 8/49 16/7 512/7

OpenStudy (anonymous):

Have you tried simplifying each expression inside the parenthesis?

OpenStudy (anonymous):

no, can I do it real quick?

OpenStudy (jdoe0001):

so.... if you were to say, as hadenthefox said, simplify say \(\bf \cfrac{8x^4x^2}{8x^7} \) what would that give you?

OpenStudy (anonymous):

96/59

OpenStudy (jdoe0001):

96/59? the "x" went MIA there

OpenStudy (anonymous):

@jdoe0001 his problem isn't using x as a variable but instead as an operator I bleieve

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

the x means multiplication & she

OpenStudy (jdoe0001):

I thought the "x" just meant "x"... hmmm so is multiplication?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so \(\bf \left( \cfrac{8\cdot 4\cdot 2}{8\cdot 7} \right)^2 \cdot \left( \cfrac{8\cdot 0}{7\cdot {-3}} \right)^3\cdot 7\cdot {-9} ?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait for the 8 x 0 it's suppouse to be 8^0 . 0 is an exponent

OpenStudy (jdoe0001):

ok so.. hmm so what's say 8 * 0 = ?

OpenStudy (anonymous):

no, the 7 is to the power of -9, not multiplied by -9

OpenStudy (jdoe0001):

hmmm lemme redo that then.. there are a few exponents I notice

OpenStudy (anonymous):

the 0, -3, & -9 are exponents

OpenStudy (jdoe0001):

ok hold the mayo

OpenStudy (anonymous):

what

OpenStudy (jdoe0001):

keep in mind this -> \(\bf a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}} \\ \quad \\ % negative exponential denominator a^{{\color{red} n}} \implies \cfrac{1}{a^{-\color{red} n}} \qquad \qquad \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}} \)

OpenStudy (anonymous):

So @TheHangingTree here are some important exponent rules to recall for this problem and